OpenStudy (anonymous):

Match the function (f) with its inverse (f-1). f(x) = x+6/2 f(x)= 1/3x + 4 f(x)=3x + 3 f(x)=7*8-4/8 f(x)= 3x + 1 Please help me!

OpenStudy (anonymous):

Pardon me, which is the original function you are talking about?

OpenStudy (anonymous):

They are all original functions..

OpenStudy (anonymous):

I need the inverse for them all

OpenStudy (anonymous):

oh, I understand, I thought this is a list of solutions where you can pick one, so if I understand this question correct - excuse me, but nowadays I have to double check - you have to find the inverse function to each of them?

OpenStudy (anonymous):

Ok, so lets do one together then.

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

The first one, doesn't really matter which one we do, they are all linear and therefore bijective. \[y=x+\frac{6}{2} \] First you want to solve for x and represent it as a function of y. \[ x = \frac{2y-6}{2} =y-3\]

OpenStudy (anonymous):

Did I read the first equation correct? or is it everything divided by two?

OpenStudy (anonymous):

\[y= \frac{x+6}{2} \neq x+3 \] (-:

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

No sorry its all divided by 2

OpenStudy (anonymous):

I said the inverse is 2x-6 is that correct?

OpenStudy (anonymous):

good so again, but same procedure, you always do it the same way, solve for x. So you will get \[ x= 2y-6 \]

OpenStudy (anonymous):

Ok good

OpenStudy (anonymous):

Exactly, at this step you just 'switch' the variables, so you represent the inverse function as a function of x again rather then a function of y and you will get \[ y=2x-6\]

OpenStudy (anonymous):

and that's what you will have to do for all of them, good exercise anyhow.

OpenStudy (anonymous):

Im confused on the next one

OpenStudy (anonymous):

is the x in the denominator? or in the numerator?

OpenStudy (anonymous):

Its being multiplied

OpenStudy (anonymous):

\[y = \frac{1}{3} x+4\] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok this means it's in the numerator. \[ y = \frac{1}{3}x+4 = \frac{x}{3}+4\] rules of multiplication

OpenStudy (anonymous):

I said its 3x-12

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so would the third one be the same, 3x-12 also?

OpenStudy (anonymous):

\[ y=3x+3\] \[ x= \frac{1}{3} (y-3) \]

OpenStudy (anonymous):

\[y= \frac{1}{3}(x-3)\]

OpenStudy (anonymous):

so what does that mean?

OpenStudy (anonymous):

that means that the inverse of the third function is what I just posted.

OpenStudy (anonymous):

you asked if it's 3x-12 also, it is not.

OpenStudy (anonymous):

Functions and inverses are something unique in that case, so different functions never share the same inverse.

OpenStudy (anonymous):

ok... is the fourth?

OpenStudy (anonymous):

the forth looks like you are missing an x term, if not then it is only a number, therefore a constant function, a constant has no inverse.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you're welcome